# GOOD: Towards Domain Generalized Orientated Object Detection

### v1

@@ Abstract
Oriented object detection has been rapidly developed in the past few years, but most of these methods assume the training and testing images are under the same statistical distribution, which is far from reality. In this paper, we propose the task of domain generalized oriented object detection, which intends to explore the generalization of oriented object detectors on arbitrary unseen target domains. Learning domain generalized oriented object detectors is particularly challenging, as the cross-domain style variation not only negatively impacts the content representation, but also leads to unreliable orientation predictions. To address these challenges, we propose a generalized oriented object detector (GOOD). After style hallucination by the emerging contrastive language-image pre-training (CLIP), it consists of two key components, namely, rotation-aware content consistency learning (RAC) and style consistency learning (SEC). The proposed RAC allows the oriented object detector to learn stable orientation representation from style-diversified samples. The proposed SEC further stabilizes the generalization ability of content representation from different image styles. Extensive experiments on multiple cross-domain settings show the state-of-the-art performance of GOOD. Source code will be publicly available.
@@ 1 Introduction
Oriented object detection occupies a unique intersection between computer vision and remote sensing [47, 12, 51, 53], involving the prediction of oriented bounding boxes for aerial objects. Given the bird’s-eye view perspective of an image sensor [56, 55, 3, 4, 2, 62], aerial objects pose arbitrary orientations. Therefore, in addition to the conventional task of accurately identifying and localizing objects [40, 30, 27], an oriented object detector must also provide precise predictions of an object’s rotation angle [35, 49, 50, 54].
@@
Existing oriented object detectors typically rely on the assumption that the source domain during training and the target domain during inference share the same statistical distribution. However, this assumption often diverges from reality. As shown in Fig. 1a, aerial images from different domains have great style variation, caused by illumination, imaging sensor, landscape and etc. A pre-trained oriented detector on Google Earth images from American landscape (e.g., SODA [8] as source domain) need to conduct inference on numerous unseen target domains, e.g., Gaofen-2 images from Chinese landscape in DOTA [47] domain (Fig. 1b). Developing a domain generalized oriented object detector enables the inference across diverse, previously unobserved target domains. Regrettably, as far as our knowledge extends, this area remains largely unexplored.
@@
Figure 1: (a) Oriented object images different domains (e.g., SODA [8], DOTA [47], FAIR1M [39], HRSC [31]) can have dramatic style variation (e.g., spatial resolution, contrast, illumination), which leads to different feature distribution; (b) Domain generalized oriented object detection aims to learn an oriented object detector from only source domain that can be well generalized to arbitrary unseen target domains.
@@
In this paper, we make an early exploration to learn domain generalized oriented object detectors. The great style variation between the source domain and unseen target domains is usually reflected from the image appearances that have different color distribution (e.g., intensity, contrast) [22, 9, 33, 5], which are the input of an oriented object detector pre-trained on a certain source domain. The different color distribution makes the pre-trained oriented object detector difficult to perceive the activation patterns and to build the feature representation as learnt on the source domain [21, 9]. The shift of activation patterns on the target domains inevitably poses two key challenges for oriented object detectors.
First of all, the shift of activation patterns accumulates from shallow to deep of an oriented object detector, and inevitably confuses and degrades the high-level representation, e.g., region of interests (RoIs). Consequently, as illustrated in Fig. 2a, the category-level representation, also known as a kind of content information, can be less reliable for a detector to precisely identify an object. A same object category from different styles can be difficult to cluster into the same category. Secondly, such statistic variation from the target domain can also lead to less representative orientation information (in Fig. 2b). This is because the orientation is usually regressed from the RoIs. As the RoIs degrade on the target domains, the orientation inevitably becomes less precise. As illustrated in Fig. 2c, the orientation of the prediction can be quite different from the ground truth.
@@
Figure 2: Two key challenges when an oriented object detector generalizes to unseen target domains. (a) degraded category prediction: multiple ships are not detected; (b) imprecise oriented prediction: the rotation angle does not align with the object orientation. (c) Analysis from feature space for both challenges. Oriented object detector: ReDet [16]; Source domain: DOTA [47]; Unseen target domain: FAIR1M [39] & HRSC [31].
@@
We propose a domain generalized oriented object detector, dubbed as GOOD, to address the above challenges. First, we aim to allow the GOOD to perceive as many as styles when training on the source domain. The emerging contrastive language-image pre-training (CLIP) [36] rests in abundant styles from web-wide images, and provides us a feasible solution. Afterwards, the proposed GOOD consists of two key components, namely, rotation-aware content consistency learning (RAC) and style consistency learning (SEC). The proposed RAC constraints both horizontal region of interests (HRoIs) and rotated region of interests (RRoIs) between the original images and style-hallucinated images, so that both the orientation and content representation is more robust to the cross-domain style variation. On the other hand, the proposed SEC poses a category-wise constraint between the original and style-hallucinated images, which allows the style-hallucinated images to perceive the same content. In this way, the classification head yields more discriminative category prediction on the unseen target domains.
@@
Extensive domain generalization settings on existing oriented object detection benchmarks (i.e., FAIR1M [39], DOTA [47], SODA [8], HRSC [31]) have been designed to benchmark domain generalized oriented object detection. Following prior driving-scene object detection works [46, 29, 42], the same object categories among each dataset are used for experiments. Specifically, four cross-domain settings, namely, FAIR1M-10 classes, SODA-7classes, DOTA-4classes and HRSC-ship, are designed. The proposed GOOD shows substantial improvement on the multiple baselines on all the generalization settings.
@@
  • We make an early exploration to the task of learning domain generalized oriented object detectors. A generalized oriented object detector, dubbed as GOOD, is proposed.
  • With CLIP-driven style hallucination, we propose a rotation-aware consistency (RAC) and a style consistency (SEC) learning to represent generalized oriented objects.
@@ 2 Related Work
Oriented Object Detection can be categorized as two-stage and one-stage. Typical works of two-stage oriented object detectors include RoI Transformer [11], ReDet [16], Oriented R-CNN [48] and etc. One-stage oriented object detectors can be further categorized into three types. The first type (e.g.,R3Det [51], $S^{2}$-Net [15]) focus on feature alignment between the regression branch and classification branch. The second type focus on angle loss regression [35, 49, 50, 53, 54]. The third type follows the RepPoint paradigm, which leverages the point prediction to bound an oriented object [19, 26]. To learn domain generalized oriented object detectors, we choose the two-stage paradigm not only because it provides a more precise description between orientation and region of interests, but also it provides more access to enrich the style and constraint the content consistency from shallow to deep.
@@
Domain Generalized Object Detection is more challenging than domain adaptive object detection [28, 25, 60]. It aims to learn an object detector from only the source domain data that can be generalized to unseen target domains [29]. More recent works focus on learning domain generalized object detectors on scenarios such as driving scenes [46, 29, 42], 3D point cloud [23, 45], UVA image [43] and etc. On the other hand, domain generalization has received increasingly attention. Multiple advanced machine learning methods (e.g., entropy regularization [61], common-specific low-rank decomposition [34], casual matching [32], extrinsic-intrinsic interaction [44], balance invariance [6], batch normalization embeddings [38]) have been proposed to enhance the generalization ability. However, to the best of our knowledge, none of these works focus on learning oriented object detectors.
@@
Vision-language Model Driven Detection Contrastive language-image pre-training (CLIP) [36] has recently become a domaint visual representation learning trend, which utilizes the image-text pairs to learn visual-semantic representation by pre-training [36, 10, 13]. For object detection community, CLIP has significantly improved the visual representation for open vocabulary object detectors [57, 14, 1, 24, 58], which involve detecting objects based on arbitrary textual descriptions. More recently, Single-DGOD [42] utilizes CLIP for domain generalized detectors on driving scenes. However, in this work, CLIP is only used for semantic augmentation, and does not need to handle the varied orientation in the oriented object detection task.
@@ 3 Methodology
Figure 3: schema on generating style-hallucinated samples by the proposed CLIP-driven style hallucination.
Figure 4: Framework overview of the proposed GOOD. After CLIP-driven style hallucination (Sec. 3.2), two key components, namely, rotation-aware consistency learning (RAC, in Sec. 3.3) and style consistency (SEC, in Sec. 3.4) learning, are involved. RAC implements content consistency on both horizontal and rotated region of interests (HRoI and RRoI), as presented in Eq. 6 and Eq. 8, respectively. SEC implements style consistency on category-wise representation from both original and style-hallucinated images (in Eq. 10).
@@ 3.1 Problem Setup & Framework Overview
Given a source domain $\mathcal{S}$ and multiple unseen target domains $\mathcal{T}_{1}$, $\cdots$, $\mathcal{T}_{k}$, $\cdots$. In general, the source domain and unseen target domains share the same label space but are under different distributions. For a certain domain $\mathcal{S}$, the joint image and label pair is denoted as $\{\mathbf{I}_{n}^{(\mathcal{S})},(x_{n}^{(\mathcal{S})},y_{n}^{(\mathcal{S})},w_{n}^{(\mathcal{S})},h_{n}^{(\mathcal{S})},\theta_{n}^{(\mathcal{S})})\}_{n=1}^{N_{s}}$, where $n=1,2,\cdots,N_{s}$. Here $x_{n}^{(\mathcal{S})},y_{n}^{(\mathcal{S})},w_{n}^{(\mathcal{S})},h_{n}^{(\mathcal{S})},\theta_{n}^{(\mathcal{S})}$ denote the upper left x-coordinate, upper left y-coordinate, width, height and oriented angle of a oriented region of interest (RRoI), which jointly describe its 2D location in image $\mathbf{I}_{n}^{(\mathcal{S})}$. Also, $N_{s}$ denotes the sample number in domain $\mathcal{S}$.
@@
Fig. 4 gives an overview of the proposed generalized oriented object detector, dubbed as GOOD. After the feature extraction from a backbone, it has two key components, namely, Rotation-aware Consistency Learning (RAC) and Style Consistency Learning (SEC), which are used to decouple the negative impact of different styles on rotation prediction and category prediction, respectively. Overall, it is backbone independent, and can be deployed into off-the-shelf two-stage oriented object detectors that has a step of rotation region of interest (RROI) representation and regression & classification prediction.

### v2

@@ Abstract
Oriented object detection has been rapidly developed in the past few years, but most of these methods assume the training and testing images are under the same statistical distribution, which is far from reality. In this paper, we propose the task of domain generalized oriented object detection, which intends to explore the generalization of oriented object detectors on arbitrary unseen target domains. Learning domain generalized oriented object detectors is particularly challenging, as the cross-domain style variation not only negatively impacts the content representation, but also leads to unreliable orientation predictions. To address these challenges, we propose a generalized oriented object detector (GOOD). After style hallucination by the emerging contrastive language-image pre-training (CLIP), it consists of two key components, namely, rotation-aware content consistency learning (RAC) and style consistency learning (SEC). The proposed RAC allows the oriented object detector to learn stable orientation representation from style-diversified samples. The proposed SEC further stabilizes the generalization ability of content representation from different image styles. Notably, both learning objectives are simple, straight-forward and easy-to-implement. Extensive experiments on multiple cross-domain settings show the state-of-the-art performance of GOOD. Source code will be publicly available.
keywords:
[…]
{highlights}
We make an early exploration to the task of learning domain generalized oriented object detectors. The evaluation protocols of this task are accordingly established.
A generalized oriented object detector, dubbed as GOOD, is proposed. With CLIP-driven style hallucination, it consists of a rotation-aware consistency (RAC) and a style consistency (SEC) learning to represent generalized oriented objects.
The proposed RAC is innovative to constrain the both horizontal and rotated regions of interest (HRoIs and RRoIs). Technically, the pre- and post- hallucinated HRoIs and RRoIs that have the most similar orientation information is aligned, while the rest are pushed as apart as possible.
@@ 1.1 Problem Statement
Oriented object detection occupies a unique intersection between computer vision and remote sensing Xia et al. (2018); Ding et al. (2021); Yang et al. (2021b, d), involving the prediction of oriented bounding boxes for objects in aerial images. Given the bird’s-eye view perspective of an image sensor Yao et al. (2022); Yang et al. (2023); Bi et al. (2021, 2022, 2020); Zhou et al. (2021), objects in aerial images pose arbitrary orientations. Therefore, in addition to the conventional task of accurately identifying and localizing objects Tian et al. (2019); Lin et al. (2017); Li et al. (2022d), an oriented object detector must also provide precise predictions of an object’s rotation angle Qian et al. (2021); Yang and Yan (2020); Yang et al. (2021a, 2022).
@@
Existing oriented object detectors typically rely on the assumption that the source domain during training and the target domain during inference share the same statistical distribution. However, this assumption often diverges from reality. As shown in Fig. 1a, aerial images from different domains have great style variation, caused by illumination, imaging sensor, geo-location, landscape and etc. Iizuka et al. (2023); Zhu et al. (2023a). A pre-trained oriented object detector on WorldView-2 images from American landscape (e.g., SODA Cheng et al. (2023) as source domain) need to conduct inference on numerous unseen target domains, e.g., Gaofen-2 images from Chinese landscape in DOTA Xia et al. (2018) domain.
@@
Figure 1: (a) Domain generalized oriented object detection aims to learn an oriented object detector from only source domain that can be well generalized to (b) arbitrary unseen target domains.
To better understand the domain shift caused by styles, we quantify the style of aerial images from different domain following the existing definition of styles Gatys et al. (2016); Huang and Belongie (2017), where the channel-wise mean and standard deviation is used. We compute these two metrics of each sample in each domain, and visualize the style distribution of these domains. As shown in Fig. 2, the style distribution of each domain is quite different, which accounts for the domain gap. Developing a domain generalized oriented object detector enables the inference across diverse and previously unobserved target domains (Fig. 1b). Regrettably, as far as our knowledge extends, this area remains largely unexplored.
Figure 2: Quantifying the styles from different aerial image domains. The per-sample style is quantified by two variables, namely, mean and standard deviation Huang and Belongie (2017). Mean demonstrates how the samples are spread out in the feature space, whereas standard deviation demonstrates how the per-sample feature varies. For each aerial image domain, the per-sample mean and standard deviation is computed and visualized. Styles from different domains show a dramatic discrepancy.
1.2 Motivation & Objective
@@
In this paper, we make an early exploration to learn domain generalized oriented object detectors. The great style variation between the source domain and unseen target domains is usually reflected from the image appearances that have different color distribution (e.g., intensity, contrast) Lee et al. (2022a); Choi et al. (2021); Nam et al. (2021); Bi et al. (2024a), which are the input of an oriented object detector pre-trained on a certain source domain. The different color distribution makes the pre-trained oriented object detector difficult to perceive the activation patterns and to build the feature representation as learned on the source domain Huang et al. (2023); Choi et al. (2021). The shift of activation patterns on the target domains inevitably poses two key challenges for oriented object detectors.
@@
Figure 3: Two key challenges when an oriented object detector generalizes to unseen target domains. (a) degraded category prediction: multiple ships are not detected. In the feature space, samples from each category are not properly clustered due to cross-domain style variation; (b) imprecise oriented prediction: the rotation angle does not align with the object orientation, which is reflected in both image space and blue arrows in feature space.
First of all, the shift of activation patterns accumulates from shallow to deep of an oriented object detector, and inevitably confuses and degrades the high-level representation, e.g., region of interests (RoIs). Consequently, as illustrated in Fig. 3a, the category-level representation, also known as a kind of content information, can be less reliable for a detector to precisely identify an object. A same object category from different styles can be difficult to cluster into the same category.
Secondly, such statistic variation from the target domain poses a unique challenge to predict rotated bounding boxes than conventional horizontal bounding boxes, i.e., less representative orientation information (in Fig. 3b). This is because the orientation is usually regressed from the RoIs. As the RoIs degrade on the target domains, the orientation inevitably becomes less precise. As illustrated in Fig. 3c, the orientation of the prediction can be quite different from the ground truth.
@@
We propose a domain generalized oriented object detector, dubbed as GOOD, to address the above challenges. First, we aim to allow the GOOD to perceive more diverse styles when training on the source domain. The emerging contrastive language-image pre-training (CLIP) Radford et al. (2021) rests in abundant styles from web-wide images, and provides us a feasible solution. Afterwards, the proposed GOOD consists of two key components, namely, rotation-aware content consistency learning (RAC) and style consistency learning (SEC). The proposed RAC constrains both horizontal region of interests (HRoIs) and rotated region of interests (RRoIs) between the original images and style-hallucinated images, so that both the orientation and content representation is more robust to the cross-domain style variation. On the other hand, the proposed SEC poses a category-wise constraint between the original and style-hallucinated images, which allows the style-hallucinated images to perceive the same content. In this way, the classification head yields more discriminative category prediction on the unseen target domains.
Notably, although both RAC and SEC are simple, straight-forward and easy-to-implement, they still advance the contrastive learning and consistency learning in multiple predominant perspectives. In contrast to the generic contrastive learning works, RAC focuses on the simultaneous constraint of both horizontal and rotated regions of interests (HRoIs and RRoIs). For both HRoIs and RRoIs, the pre- and post- hallucinated regions of interests that have the most similar orientation information are aligned, while the rest are pushed as apart as possible. It allows the orientation information on unseen target domains to be more stable. Besides, though the Jensen-Shannon Divergence is commonly used for consistency constraint, the proposed SEC is innovative in two-fold. Firstly, it brings up a new style hallucination method from per-layer vision-language foundation model (VFM) features. In contrast, existing style hallucination methods usually only hallucinate the last-layer feature from pre-trained ImageNet encoder. Our method allows a more fine-grained and progressive hallucination from CLIP encoder with more diverse styles. Besides, hallucinating from CLIP features has also not been explored before.
@@
Last, but not least, it is highly important to establish the evaluation protocols for domain generalized oriented object detection, which provides a solid baseline for future study. Extensive domain generalization settings on existing oriented object detection benchmarks (i.e., , FAIR1M Sun et al. (2022), DOTA Xia et al. (2018), SODA Cheng et al. (2023), HRSC Liu et al. (2016)) have been designed to benchmark domain generalized oriented object detection. Following prior driving-scene object detection works Wu and Deng (2022); Lin et al. (2021); Vidit et al. (2023), the common object categories among each dataset are used for experiments. Specifically, four cross-domain settings, namely, FAIR1M-10 classes, SODA-7 classes, DOTA-4 classes and HRSC-ship, are involved. The proposed GOOD shows substantial improvement on the multiple baselines on all the generalization settings.
1.3 Contribution
@@
  • We make an early exploration to the task of learning domain generalized oriented object detectors. The evaluation protocols of this task are accordingly established.
  • A generalized oriented object detector, dubbed as GOOD, is proposed. With CLIP-driven style hallucination, it consists of a rotation-aware consistency (RAC) and a style consistency (SEC) learning to represent generalized oriented objects.
  • The proposed RAC is innovative to constrain the both horizontal and rotated regions of interest (HRoIs and RRoIs). Technically, the pre- and post- hallucinated HRoIs and RRoIs that have the most similar orientation information is aligned, while the rest are pushed as apart as possible.
@@
The remainder of this paper is organized as follows. In Section 2, related work is provided. In Section 3, the proposed method is demonstrated. In Section 4, we report and discuss the experiments on three aerial image scene classification benchmarks. Finally in Section 5, the conclusion is drawn.
@@ 2 Related Work
2.1 Oriented Object Detection
Existing works can be categorized into two types, namely, two-stage and one-stage. Horizontal bounding box Deng et al. (2018); Liu et al. (2021); Zhu et al. (2023b); Tan et al. (2023) is less precise to depict the spatial position and the orientation of an object. Early rotated bounding box supervised works usually modified the anchor related modules Tian et al. (2019); Lin et al. (2017) for oriented objects.
@@ 2.1 Oriented Object Detection
Typical works of two-stage oriented object detectors include RoI Transformer Ding et al. (2019), ReDet Han et al. (2021b), Oriented R-CNN Xie et al. (2021), JDet Xiao et al. (2024) and etc. . One-stage oriented object detectors can be further categorized into three types. The first type (e.g., R3Det Yang et al. (2021b), $S^{2}$-Net Han et al. (2021a), DCFL Xu et al. (2023)) focus on feature alignment between the regression branch and classification branch. The second type focus on angle loss regression Qian et al. (2021); Yang and Yan (2020); Yang et al. (2021a, d, 2022); Yu and Da (2023); Xu et al. (2024). The third type follows the RepPoint paradigm, which leverages the point prediction to bound an oriented object Hou et al. (2023); Li et al. (2022c).
More recently, some recent works jointly leverage the advantages of both the first paradigm and the second paradigm. For example, Qiao et al. (2023) proposes an adaptive angle optimization head on top of the conventional regional proposal branch. Ming et al. (2023) proposes a framework that jointly improves the image representation from the region proposal network, the orientation estimation branch and the label assignment branch.
However, these oriented object detectors have an obvious weakness that they assume the training and inference objects are under the same distribution. When there is large feature distribution shift caused by style variation between the source and target domains, both the rotation angle prediction and the category prediction can suffer severe degradation. Several recent works tackle this challenge by domain adaptation, which still need both source domain and target domain when training Liu et al. (2024a, b). Ideally, the oriented object detector, only trained on a source domain, is supposed to be well generalized to arbitrary unseen target domains.
To learn domain generalized oriented object detectors, we choose the two-stage paradigm not only because it provides a more precise description between orientation and region of interests, but also it provides more access to enrich the style and constrain the content consistency from shallow to deep.
2.2 Domain Generalized Object Detection
@@
Domain generalized object detection is more challenging than domain adaptive object detection Li et al. (2022e, b); Zhao and Wang (2022). It aims to learn an object detector from only the source domain data that can be generalized to unseen target domains Lin et al. (2021). More recent works focus on learning domain generalized object detectors on scenarios such as driving scenes Wu and Deng (2022); Lin et al. (2021); Vidit et al. (2023), 3D point cloud Lehner et al. (2022); Wang et al. (2023b), UVA image Wang et al. (2023a) and etc. . On the other hand, domain generalization has received increasing attention. Multiple advanced machine learning methods (e.g., entropy regularization Zhao et al. (2020), common-specific low-rank decomposition Piratla et al. (2020), casual matching Mahajan et al. (2021), extrinsic-intrinsic interaction Wang et al. (2020), balance invariance Chattopadhyay et al. (2020), batch normalization embeddings Segu et al. (2023)) have been proposed to enhance the generalization ability. An extensive survey can refer to Zhou et al. (2022). Meanwhile, some recent works studied domain generalization under the single source domain setting Qiao et al. (2020); Peng et al. (2022); Zhou et al. (2020); Bi et al. (2024b); Yi et al. (2024b, a) or unsupervised setting Harary et al. (2022); Hu and Lee (2022).
However, to the best of our knowledge, none of these works focus on learning oriented object detectors. They predict horizontal bounding boxes Wu and Deng (2022); Liu et al. (2023); Vidit et al. (2023), but are not applicable to regress the orientation angle. In other words, it is difficult to fit them into the oriented object detection task.
2.3 Vision-language Model Driven Detection
@@
Contrastive language-image pre-training (CLIP) Radford et al. (2021) has recently become a domaint visual representation learning trend, which utilizes the image-text pairs to learn visual-semantic representation by pre-training Radford et al. (2021); Desai and Johnson (2021); Engilberge et al. (2018). For object detection community, CLIP has significantly improved the visual representation for open vocabulary object detectors Zareian et al. (2021); Gu et al. (2022); Bangalath et al. (2022); Li et al. (2022a); Zhang et al. (2022), which involve detecting objects based on arbitrary textual descriptions.
More recently, Single-DGOD Vidit et al. (2023) utilizes CLIP for domain generalized detectors on driving scenes. The general idea is that, CLIP is pre-trained on large-scale image-text pairs, which contains much more abundant style information than naive augmentation or ImageNet pre-trained model. It provides the possibility to enrich the image styles. In Vidit et al. (2023), although the proposed DomainGen does not rely on the CLIP image encoder, it extracts a more generalized feature embedding from the CLIP text encoder, given the inputted scene description and domain description. It also relies on the stronger generalization ability of the large-scale language-image pairs for pre-training. Besides, it is not able to handle the varied orientation in the oriented object detection task. Meanwhile, style augmentation is a common and effective path to enhance the generalization ability to unseen domains. Some recent works have leveraged CLIP to enhance the style diversity Cho et al. (2023); Patashnik et al. (2021); Kwon and Ye (2022), which motivate us to implement style hallucination by CLIP on source domain images.
@@ 3 Methodology
Figure 4: Framework overview of the proposed GOOD. After CLIP-driven style hallucination (Sec. 3.2), two key components, namely, rotation-aware consistency learning (RAC, in Sec. 3.3) and style consistency learning (SEC, in Sec. 3.4), are involved. RAC implements content consistency on both horizontal and rotated region of interests (HRoI and RRoI), as presented in Eq. 6 and Eq. 8, respectively. SEC implements style consistency on category-wise representation from both original and style-hallucinated images (in Eq. 10).
@@ 3.1 Problem Setup & Framework Overview
Assume we have a source domain $\mathcal{S}$ and multiple unseen target domains $\mathcal{T}_{1}$, $\cdots$, $\mathcal{T}_{k}$, $\cdots$. In general, the source domain and unseen target domains share the same label space but are under different distributions. For domain $\mathcal{S}$, the joint image and label pair is $\{\mathbf{I}_{n}^{(\mathcal{S})},(x_{n}^{(\mathcal{S})},y_{n}^{(\mathcal{S})},w_{n}^{(\mathcal{S})},h_{n}^{(\mathcal{S})},\theta_{n}^{(\mathcal{S})})\}_{n=1}^{N_{s}}$, where $n=1,2,\cdots,N_{s}$. Here $x_{n}^{(\mathcal{S})},y_{n}^{(\mathcal{S})},w_{n}^{(\mathcal{S})},h_{n}^{(\mathcal{S})},\theta_{n}^{(\mathcal{S})}$ denote the upper left x-coordinate, upper left y-coordinate, width, height and oriented angle of an oriented region of interest (RRoI), which jointly describe its 2D location in image $\mathbf{I}_{n}^{(\mathcal{S})}$. $N_{s}$ denotes the sample number in domain $\mathcal{S}$.
@@
Fig. 4 gives an overview of the proposed generalized oriented object detector, dubbed as GOOD. After the feature extraction from a backbone, it has two key components, namely, Rotation-aware Consistency Learning (RAC) and Style Consistency Learning (SEC), which are used to decouple the negative impact of different styles on rotation prediction and category prediction, respectively. Overall, it is backbone independent, and can be deployed into off-the-shelf two-stage oriented object detectors that have a step of rotation region of interest (RROI) representation and regression & classification prediction.